OpenStudy (burgershop101):

Algebraically prove that the difference of the squares of any two consecutive integers is an odd integer.

OpenStudy (3mar):

Can you prove it using an example?

OpenStudy (burgershop101):

yeh

OpenStudy (sshayer):

let x and x+1 be consecutive integers.Clearly x is an integer. \[(x+1)^2-x^2=x^2+2x+1-x^2=2x+1\] Because x is an integer therefore 2x+1 is an odd integer.

OpenStudy (3mar):

Thank you good and brief explanation, @sshayer !